# WorldCuisines: A Massive-Scale Benchmark for Multilingual and Multicultural Visual Question Answering on Global Cuisines

### v1

@@ 1 Introduction
However, there is limited research on evaluating the multicultural capabilities of VLMs, particularly in terms of multilinguality. The study by Romero et al. (2024) introduces visual question answering (VQA) from a multicultural perspective, but it mainly focuses on knowledge and situational context at a specific moment, which does not fully assess the ability of VLMs to reason and differentiate between cultures within a single question. Moreover, another study on food VQA is limited to Chinese culture and does not explore the broader spectrum of global cultures Li et al. (2024). An earlier investigation into cultural bias in language models also found that cultural knowledge is lacking Naous et al. (2023). Therefore, further research is necessary to address these limitations and enhance our understanding of VLMs’ multicultural and multilingual capabilities.
@@
  • We create multilingual templates for queries and context (i.e., question part of QA pair) and keeping language varieties (i.e., dialects and registers) by creating translations with different inflections, articles, and contractions. We aim for naturalness in each translation and use appropriate inflections for place names.
@@ 2.2.2 Metadata Annotation
Table 3: Accuracy (%) results of WC-VQA for Test Large (60k). MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively.
@@ 3.1 Experimental Setup
We evaluate our benchmark on various available VLMs, including 14 open-source models and 3 proprietary models. During the inference of the open-source model, we use 16-bit floating point and employ greedy decoding. We access the proprietary models through API. The complete list of the models is available in Table 3.
@@ 5 Related Work
Several prior studies have focused on developing culturally relevant VQA benchmarks, including FM-IQA Gao et al. (2015), MCVQA Gupta et al. (2020), xGQA Pfeiffer et al. (2022), MaXM Changpinyo et al. (2023), MTVQA Tang et al. (2024), MABL Kabra et al. (2023), MAPS Liu et al. (2024a), and MaRVL Liu et al. (2021). Additionally, CVQA Romero et al. (2024) and CulturalVQA Nayak et al. (2024) provide VQA datasets that cover various regions and diverse topics, including food, with CVQA also offering questions in multiple languages alongside English translations. SEA-VQA Urailertprasert et al. (2024) specifically benchmarks the South East Asian region. In contrast, FoodieQA Li et al. (2024) and World Wide Dishes Magomere et al. (2024) are focused exclusively on food-related benchmarks. Our work is similarly motivated by using food as a cultural proxy, but it distinguishes itself with a significantly larger dataset and broader coverage of languages and dialects.
@@
Recent advancements in VLMs have led to the emergence of multi-modal LLMs that can process both images and text. LLaVA Liu et al. (2024c) exemplifies this approach by utilizing Vicuna Zheng et al. (2023) as an image encoder, thereby enhancing visual understanding. This architecture has set a precedent for other VLMs, including Qwen2-VL Bai et al. (2023), Llama 3.2 Dubey et al. (2024), Pixtral, Phi-3.5 Vision Abdin et al. (2024), Molmo Deitke et al. (2024), and NVLM Dai et al. (2024), each leveraging their respective large language models for multi-modal tasks. In a specialized application, FoodLMM Yin et al. (2023) focuses specifically on the food domain, training on publicly available food datasets and conversational data generated by GPT-4 Achiam et al. (2023). Our work evaluates the capabilities of these models within the food domain, offering insights into their performance and potential applications in culinary-related tasks across multicultural settings.
@@ Acknowledgements
We extend our gratitude to everyone who has supported our project, particularly the numerous annotators who provided rigorous and comprehensive annotations and conducted thorough quality checks. Special thanks go to Francesca Porcu for her assistance with the Sardinian language and to Shintaro Ozaki for his help with Japanese. We also greatly appreciate the valuable contributions of Nayeon Lee and Wenliang Dai for their insightful discussions and for integrating NVLM into our benchmark.
@@ Appendix B Open-Source Collaborative Effort
The WorldCuisines data collection and benchmark construction is a fully open-source project. We invite contributions from researchers, practitioners, and grassroots communities, such as local NLP communities, who are interested in participating. Contributions can include data collection, annotation, quality checks, and evaluation. To ensure high-quality data, we engage native speakers of local languages in the annotation process with strict quality control (QC). The contributors who provide substantial contribution are invited to have co-authorship on this paper. We follow the guidelines from ACL for authorship eligibility https://www.aclweb.org/adminwiki/index.php/Authorship_Changes_Policy_for_ACL_Conference_Papers. Our goal is to develop a resource and benchmark that will have a meaningful impact on future research. To achieve this, we are dedicated to expanding language coverage and ensuring that contributions are as inclusive and diverse as possible.
@@ C.2 Secondary Metric: BERTScore
Table 7: Accuracy (%) results of WC-VQA. MCQ and OEQ indicate multiple-choice question and open-ended question, respectively.
@@
Table 8: BERTScore results of WC-VQA. Only the results from open-ended (OEQ) are used.

### v2

@@ 1 Introduction
However, there is limited research on evaluating the multicultural capabilities of VLMs, particularly in terms of multilinguality. The study by Romero et al. (2024) introduce visual question answering (VQA) from a multicultural perspective, but it mainly focuses on knowledge and situational context at a specific moment, which does not fully assess the ability of VLMs to reason and differentiate between cultures within a single question. Moreover, another study on food VQA is limited to Chinese culture and does not explore the broader spectrum of global cultures Li et al. (2024b). An earlier investigation into cultural bias in language models also found that cultural knowledge is lacking Naous et al. (2023). Therefore, further research is necessary to address these limitations and enhance our understanding of VLMs’ multicultural and multilingual capabilities.
@@
  • We create multilingual templates for queries and context (such as the questions in QA pairs) while preserving language varieties, including dialects and registers. This is achieved by creating translations that incorporate different inflections, articles, and contractions. Our goal is to ensure naturalness in each translation and to use appropriate inflections for place names.
@@ 2.2.2 Metadata Annotation
Table 3: Accuracy (%) results of WC-VQA for Test Large (60k). MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively. We employ an optimized prompt provided by the authors (see Subsection D.1 in the Appendix for further details).
@@ 3.1 Experimental Setup
We evaluate our benchmark on various available VLMs, including 15 open-source models and 3 proprietary models. During the inference of the open-source model, we use 16-bit floating point and employ greedy decoding. We access the proprietary models through API. The complete list of the models is available in Table 3.
@@ 5 Related Work
Several prior studies have focused on developing culturally relevant VQA benchmarks, including FM-IQA Gao et al. (2015), MCVQA Gupta et al. (2020), xGQA Pfeiffer et al. (2022), MaXM Changpinyo et al. (2023), MTVQA Tang et al. (2024), MABL Kabra et al. (2023), MAPS Liu et al. (2024a), and MaRVL Liu et al. (2021). Additionally, CVQA Romero et al. (2024) and CulturalVQA Nayak et al. (2024) provide VQA datasets that cover various regions and diverse topics, including food, with CVQA also offering questions in multiple languages alongside English translations. SEA-VQA Urailertprasert et al. (2024) specifically benchmarks the South East Asian region. In contrast, FoodieQA Li et al. (2024b) and World Wide Dishes Magomere et al. (2024) are focused exclusively on food-related benchmarks. Our work is similarly motivated by using food as a cultural proxy, but it distinguishes itself with a significantly larger dataset and broader coverage of languages and dialects.
@@
Recent advancements in VLMs have led to the emergence of multi-modal LLMs that can process both images and text. LLaVA Liu et al. (2024c) exemplifies this approach by utilizing Vicuna Zheng et al. (2023) as an image encoder, thereby enhancing visual understanding. This architecture has set a precedent for other VLMs, including Qwen2-VL Bai et al. (2023), Llama 3.2 Dubey et al. (2024), Pixtral Agrawal et al. (2024), Phi-3.5 Vision Abdin et al. (2024), Molmo Deitke et al. (2024), Aria Li et al. (2024a), Pangea Yue et al. (2024), and NVLM Dai et al. (2024), each leveraging their respective large language models for multi-modal tasks. In a specialized application, FoodLMM Yin et al. (2023) focuses specifically on the food domain, training on publicly available food datasets and conversational data generated by GPT-4 Achiam et al. (2023). Our work evaluates the capabilities of these models within the food domain, offering insights into their performance and potential applications in culinary-related tasks across multicultural settings.
@@ Acknowledgements
We extend our gratitude to everyone who has supported our project, especially the numerous annotators who provided meticulous and comprehensive annotations and conducted thorough quality checks. Special thanks to Francesca Porcu for her assistance with the Sardinian language and to Shintaro Ozaki for his help with Japanese. We are also deeply appreciative of Nayeon Lee and Wenliang Dai for their insightful discussions and for integrating NVLM into our benchmark. Additionally, we thank Xiang Yue and Yueqi Song for their help in integrating Pangea into our benchmark.
@@ Appendix B Open-Source Collaborative Effort
The WorldCuisines data collection and benchmark construction is a fully open-source project. We invite contributions from researchers, practitioners, and grassroots communities, such as local NLP communities, who are interested in participating. Contributions can include data collection, annotation, quality checks, and evaluation. To ensure high-quality data, we engage native speakers of local languages in the annotation process with strict quality control (QC). The contributors who provide substantial contribution are invited to have co-authorship on this paper. We follow the guidelines from ACL for authorship eligibility. [^6] The ACL guidelines can be found at https://www.aclweb.org/adminwiki/index.php/Authorship_Changes_Policy_for_ACL_Conference_Papers. Our goal is to develop a resource and benchmark that will have a meaningful impact on future research. To achieve this, we are dedicated to expanding language coverage and ensuring that contributions are as inclusive and diverse as possible.
@@ C.2 Secondary Metric: BERTScore
Table 7: Accuracy (%) results of WC-VQA. MCQ and OEQ indicate multiple-choice question and open-ended question, respectively. Best and second-best are bolded and underlined, respectively. We employ an optimized prompt provided by the authors (see Subsection D.1 in the Appendix for further details).
@@
Table 8: BERTScore results of WC-VQA. Only the results from open-ended (OEQ) are used. Best and second-best are bolded and underlined, respectively.
@@
Appendix D Evaluation
D.1 Prompt Sensitivity
We use the same prompts for all models, with the exception of the Pangea 7B model Yue et al. (2024). This model is particularly sensitive and lacks robustness in handling diverse prompt instructions, often struggling to follow instructions accurately, especially in multiple-choice questions (MCQs), unless a specific template is applied. In contrast, models like Llama 3.2 Instruct and Qwen2 VL Instruct are more adaptable to varied instructions. After consulting with the authors, we adopted the prompt “Answer with the option letter from the given choices directly.” for MCQ queries when using the Pangea 7B model.